Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 8:24 Saturday 09 December 2023 (62685) Print this report
Operator Report - Initial Status

Quick Summary: ISYS in standard state; All Suspensions Loops closed; EIB act0 close to saturation, all other SBE loops close and in stanrda condition; INJ LaserChiller Laser Flow decreasing; TCS NI Chiller flow decreasing;
Activities ongoing since this morning: Daily Calibrations;

Images attached to this report
Comments to this report:
menzione - 22:43 Sunday 10 December 2023 (62694) Print this report

concerning the TCS NI Chiller flow decreasing, Yesterday at 14:00 UTC I went in TCS chiller room to check the status of the NI aux pipe, I noticed that it was a bit strangled. I stretched it out in a correct position and the flow came back to the normal value. (plot attached)

Images attached to this comment
Search Help